# Autoregressive Ranking: Bridging the Gap Between Dual and Cross Encoders

### v1

@@ Abstract
Dual and cross encoders have long been mainstays of information retrieval (IR), but are being challenged by the emergent capabilities of LLMs. An LLM-based approach we term pointwise generative ranking — generating tokens the length of a single docID as opposed to a list in order to enable ranking via beam search — combines efficiency and expressivity benefits while leveraging the in-context capabilities of Causal Transformers. Although there is ample evidence to suggest that pretrained LLMs are well-suited for ranking, we find that the vast majority of LLM-based approaches rely on next-token prediction, a loss function which is fundamentally rank-agnostic (and especially so with pointwise supervision). In this paper, we first prove that the expressivity of pointwise generative ranking with multi-token docIDs is superior to that of dual encoders. We then propose SToICaL — a Simple Token-Item Calibrated Loss — which can incorporate rank-aware supervision at both the item and token levels within the pointwise setup. We run a suite of experiments on ranking tasks derived from WordNet (fellbaum1998wordnet) and ESCI (reddy2022shopping). Two variants of SToICaL successfully suppress the probability of invalid docID generations and improve on common ranking metrics beyond top-1 retrieval.
(a) Dual Encoder (DE)
(b) Cross Encoder (CE)
(c) Autoregressive Ranking (ARR)
@@ 1 Introduction
The task of ranking documents by their relevance to a user query is a cornerstone of modern information retrieval. The dominant paradigm for this task is a multi-stage pipeline. The first stage, retrieval, aims to efficiently select a candidate set of documents from a massive corpus. This is typically accomplished using Dual Encoder (DE) models, which represent queries and documents as dense vectors and use fast approximate nearest neighbor (ANN) search to find the top matches (karpukhin2020dense; guo2020accelerating). While efficient, the expressive power of DEs is inherently limited, as the rich interaction between a query and a document is compressed into a single vector similarity. Consequently, a second reranking stage is often required. In this stage, a more powerful but computationally expensive Cross Encoder (CE) model jointly processes the query and each candidate document to produce a more accurate relevance score (nogueira2019passage). CEs are infeasible for first-stage retrieval due to their prohibitive cost, which scales linearly with the corpus size.
@@
The recent success of large language models (LLMs) has inspired a new, unified approach to retrieval and ranking in lieu of the two-stage approach (metzler2021rethinking; li2023large; lee2024can). In this paradigm, an LLM is trained to function as a generative retrieval model, autoregressively decoding the relevant documents or their identifiers (docIDs) token-by-token, conditioned on the query (tay2022transformer). For ranking, this model can generate a sorted list of documents by employing beam search, effectively ranking documents by their conditional generation probability. This approach, which we refer to as Autoregressive Ranking (ARR) (see Figure˜1 for an illustration), offers several compelling advantages over the traditional DE/CE pipeline. [^2] The theoretical focus of this work on the ranking stage of IR, as well as our proposed loss function which presupposes token-by-token generation in decoder LLMs, justify the narrower term “autoregressive ranking” over the broader “generative retrieval” in the context of this exposition. It collapses the two-stage process into a single model, eliminating the need for a separate ANN index. Furthermore, it is more efficient than CEs as it does not require scoring every document individually, while preserving CEs’ cross-attention over already generated docID tokens. Crucially, it has also been argued that by generating documents token-by-token, ARR models can capture deeper and more nuanced query-document interactions than DEs (yuan2024generative); while intuitive, the evidence for this has been primarily empirical.
Paper Contributions.
This paper addresses two critical gaps in the understanding and optimization of ARR models for ranking.
@@
1) We provide a theoretical foundation for the superior expressive capability of ARR over DEs. We conduct a rigorous analysis of the embedding geometry required for ranking. We prove that for a DE to correctly rank any permutation of $k$ documents, its embedding dimension must grow linearly with $k$. In contrast, we show that an ARR model with a constant hidden dimension is theoretically sufficient to rank an arbitrary number of documents. This result offers a formal explanation for the empirical advantages of generative approaches.
@@
2) We introduce SToICaL, a rank-aware generalization of next-token prediction loss for enhanced ranking capabilities. Most existing generative retrieval methods are trained with standard next-token prediction loss, an objective which fails to explicitly model the relative ordering of documents, which is crucial for ranking. To address this, we introduce a new loss function designed to make the LLM aware of the entire ranked list during training, while keeping the generation and signal length on the order of a single docID rather than a full list, enabling ranking via beam search.
@@ 2 Related Work
Types of Losses for Retrieval and Ranking
Training objectives for retrieval and ranking can be broadly grouped into the categories of pointwise, pairwise, and listwise (liu2009learning). A pointwise loss of the form $\mathcal{L}(q,d)$ encourages the model to assign high scores to more relevant items and vice versa; each training example consists of a single document to penalize independently of the rest (e.g., see chen2009ranking for a classical pointwise loss formulation). Pairwise loss functions (cao2007learning) of the form $\mathcal{L}(q,d^{+},d^{-})$ strive to maximize the margin between relevant and irrelevant documents, and commonly take the form of triplet losses (positive score, negative score, and a margin). Listwise losses of the form $\mathcal{L}(q,d_{1},\ldots,d_{n_{q}})$ jointly take as input the model scores of all documents under consideration for a given query, and can encourage the most flexible scoring configurations at the granularity of permutations rather than individual items. In this context, the Plackett-Luce probability model over permutations (luce1959individual; plackett1975analysis) forms the basis for a number of listwise losses, including ListMLE (xia2008listwise). The “Binary”, “Pairwise”, and “Coupled” variants of the RankDistil (reddi2021rankdistil) framework for top-$k$ ranking reflect the utility of each of the three types of losses.
@@
Autoregressive decoder LLMs introduce the facet of text generation on top of the simple API of “model-as-document-scorer”. The score-based loss formulations discussed above remain applicable insofar as a docID’s log-probability can be obtained by summing the log-probabilities of its constituent tokens; notable LLM-based reranking methods such as UPR (sachan2022improving) and PaRaDe (drozdov2023parade) rely specifically on LLM-based log-likelihood to obtain scores. At the same time, the terminology of “pointwise” vs “listwise” naturally acquires a second meaning, i.e., whether the text string which the LLM generates comprises a single docID or a ranked list of docIDs (separated by a delimiter and in descending order, as per common convention). Our “pointwise” approach relates to both these senses (i.e., scoring just one document, or generating just a single docID string), since steering LLM log-probabilities to be more rank-aware also directly impacts the rankings produced by a decoding strategy like beam search at inference time. The generative approach naturally allows us to perform ranking with a single model call, in contrast to CEs which must evaluate the score for every document separately. Meanwhile, the compact pointwise output format mitigates snowballing errors and is efficient to train under our rank-aware modification to next-token prediction.
@@
A handful of examples illustrate the diverse reranking paradigms afforded by generative LLMs. RankVicuna (pradeep2023rankvicuna) and RankZephyr (pradeep2023rankzephyr) explore training/distillation choices, using instruction tuning with next token prediction, to achieve a robust open-source zero-shot LLM for listwise ranking. LTRGR (li2024learning), on top of a next-token-prediction-based “learning-to-generate” phase, introduces a “learning-to-rank” phase that maximizes the margin between the autoregressively-computed scores of relevant and irrelevant passage identifiers. RecRanker (luo2024recranker) trains an LLM with instruction tuning on a combination of pointwise, pairwise, and listwise ranking objectives for top-$k$ sequential recommendation, and at inference time aggregates the three types of responses into a hybrid ranking by means of respective utility functions. FIRST (reddy2024first) uses the logits for the very first generated token to accelerate inference for models trained on listwise ranking. qin2023large introduce Pairwise Ranking Prompting (PRP), making pairwise comparisons that require post-hoc aggregation. tang2023found propose Permutation Self-Consistency as a robust listwise ranking strategy to mitigates position bias. GenRank (huang2025towards) identify the autoregressive generation mechanism as critical to the effectiveness of generative ranking, more so than the pre-training paradigm. sun2023chatgpt explore the ranking capabilities of ChatGPT and GPT-4 and propose a distillation technique to learn smaller specialized rankers. gupta2025scalable introduce BlockRank, efficiently tailoring the attention mechanism to in-context ranking. We argue that autoregressive methods which employ next-token prediction or instruction fine-tuning are relying on a token-level loss which is fundamentally not rank-aware. Our proposed loss function is a simple, principled and effective injection of rank-awareness into next-token prediction — a loss function which is ubiquitous to autoregressive models.
@@
Differentiable Search Index (DSI) (tay2022transformer) aims to instill the docID index in model weights as part of the training objective. A critical component of parametric generative retrieval is effective docID design (e.g., li2023multiview) [^3] For a literature review on docID design see Section 3 of https://ecir2024-generativeir.github.io/.; we consider the question of docID learning for the ESCI dataset in Section˜5.1. Note that we opt to use in-context ranking (gupta2025scalable), which leverages the LLM’s in-context reasoning and copying abilities to perform retrieval and ranking from a pool of documents, with their docIDs provided as part of the prompt (cf. the “Large Search Model” paradigm of wang2024large). In the case of in-context ranking, one can interpret the longer prompt as an “augmented query”, illustrated by the red “query” dots of Figure˜1(c) — the figure is intended to generalize the DSI with the in-context ranking approach. Our theory (Section˜3) and loss function (Section˜4) similarly generalizes to parametric as well as in-context ranking.
@@ 3 The Capacity for Ranking
Let $\mathscr{Q}$ be a set of queries and $\mathscr{D}$ be a corpus of documents (or their docIDs).
@@
We consider a ranking task where for each query $q\in\mathscr{Q}$, the goal is to generate a list $\mathscr{L}(q)=[d_{1}(q),\ldots,d_{k}(q)]$ of relevant documents, where each $d_{i}(q)\in\mathscr{D}$. The ordering of this list is significant: for any $i<j$, $d_{i}(q)$ is considered more relevant to $q$ than $d_{j}(q)$, denoted by $d_{i}(q)\succ d_{j}(q)$.
Autoregressive Ranking (ARR).
We define Autoregressive Ranking (ARR) as the use of a causal language model for ranking tasks. Such a model predicts the probability of a sequence of tokens conditioned on its preceding context. To formalize this, let $\mathcal{V}$ be a finite vocabulary set and $\mathcal{T}:\mathscr{Q}\cup\mathscr{D}\rightarrow\mathscr{V}^{*}$ be an associated tokenizer. Here, $\mathscr{V}^{*}\coloneqq\{[v_{1},\ldots,v_{t}]\mid t\geq 1,v_{i}\in\mathscr{V}\}$ denotes the set of sequences of tokens of arbitrary lengths.
Given a query $q$, an ARR model defines a conditional probability distribution over document token sequences. The probability of any document $d\in\mathscr{D}$ conditioned on a query $q\in\mathscr{Q}$ is the product of $d$’s token probabilities, calculated autoregressively:
$p(d\,|\,q)\coloneqq P(\mathcal{T}(d)\,|\,\mathcal{T}(q))=\prod_{i=1}^{|\mathcal{T}(d)|}P(v_{i}\,|\,\mathcal{T}(q),v_{1},\ldots,v_{i-1}),$ (1)
where $\mathcal{T}(d)\coloneqq[v_{1},\ldots,v_{|\mathcal{T}(d)|}]$. Conceptually, a ranking of the documents can be readily obtained by calculating this probability for every document and sorting the results in descending order. [^4] In practice, this exhaustive scoring is computationally prohibitive for a large-scale corpus. Fortunately, the autoregressive nature of causal models offers a more efficient solution. Instead of scoring, we can directly generate the high-probability documents using greedy decoding or beam search. By employing beam search, we can generate the most likely documents conditioned on the query, directly yielding a ranked list without the need for exhaustive computation.
Scoring-Based Ranking Architecture
Generative autoregressive rankers (ARRs) as well as dual encoders (DEs) are both examples of a scoring-based ranking architecture. Such architectures work by generating a score for each document given a query $q$, with the objective that the scores, when sorted in descending order, align with the ground-truth ordering of the relevant documents, and that scores corresponding to irrelevant documents are smaller than scores for relevant ones.
@@
A (parametric) scoring-based ranking architecture is a function $f(q,d;\theta):\mathscr{Q}\times\mathscr{D}\rightarrow\mathbb{R}$, which takes a pair $q,d$ and produces a scalar score interpreted as the similarity. Here, $\theta$ denotes learnable model parameters.
In particular, we say that a scoring-based architecture $f(q,d;\theta)$ can solve a ranking task if there exists a $\theta$ such that the scores solve the ranking task for all queries in $\mathscr{Q}$.
To simplify our analysis, we consider a particular case of the ranking task defined as follows.
@@
Given $\mathscr{Q}$ and $\mathscr{D}$, we call a ranking task complete if:
  • for each $q\in\mathscr{Q}$, all documents are relevant, i.e., $\mathscr{D}(q)=\mathscr{D}$, and
  • for any ordering $\pi(\mathscr{D})$ of the documents $\mathscr{D}$, there exists a $q\in\mathscr{Q}$ such that its ranked list of relevant documents $\mathscr{L}(q)$ is the same as $\pi(\mathscr{D})$.
We proceed to show that the architectural capacity of ARR for ranking tasks is not limited in the same way as that of a DE.
@@ 3.1 Ranking via Dual Encoders
In this section, we start with an analysis of a DE’s capacity for the ranking task. Our result shows that an embedding dimension that grows linearly with the number of relevant documents is necessary for a DE to correctly rank them. Subsequently, we analyze ARR models and show that their capability to rank documents is independent of the number of documents. Such a result rigorously establishes the advantage of ARRs over DEs.
In DEs, the query and document are embedded into the same Euclidean space, with the negated Euclidean distance as the similarity score. Specifically, a dual encoder generates a similarity score for each pair of $q\in\mathscr{Q}$ and $d\in\mathscr{D}$ as
$f_{\text{DE}}(q,d;\theta)=-\|E_{Q}(q;\theta_{\mathscr{Q}})-E_{D}(d;\theta_{\mathscr{D}})\|_{2}$ (2)
where $E_{Q}:\mathscr{Q}\to\mathbb{R}^{n}$ and $E_{D}:\mathscr{D}\to\mathbb{R}^{n}$ are query and document encoders, respectively, and $n$ is the dimension of the embedding space. Here, we assume that $E_{Q}$ and $E_{D}$ can have an arbitrary architecture. This means that they can be taken to be as expressive as needed when studying whether or not DEs are expressive enough to solve the ranking task.
Our main result for DEs is the following.
Theorem 3.1 (Insufficiency of Dual Encoders for Complete Ranking).
Let $f_{\text{DE}}(q,d;\theta)$ be a dual encoder architecture defined in Equation˜2 with some embedding dimension $n$. Then there does not exist a $\theta$ such that $f_{\text{DE}}(q,d;\theta)$ solves a complete ranking task when
@@
Let $N_{n,p}(k)$ denote the maximum number of distinct distance permutations generated by $k$ sites in $\mathbb{R}^{n}$ with the $L_{p}$ metric (skala2009counting). From (skala2009counting, Corollary 8), $N_{n,2}(k)\leq k^{2n}$. When Inequality (3) holds, we have $k^{2n}<k^{!}$, and thus $n$-dimensional embeddings are unable to capture all $k!$ permutations required in a complete ranking task. ∎
Using Stirling’s formula which states $\ln(k!)=k\ln k-k+O(\ln(k))$, the condition in Equation˜3 becomes $n<\frac{k}{2}-\frac{k}{2\ln k}+O(1)$. Since $k/\ln k$ grows sublinearly in $k$, this result states that the embedding dimension $n$ of the DE needs to grow linearly with the number of documents $k$.
@@
3.2 Ranking via (Pointwise) Generative Rankers
We turn our attention to the ability of an ARR model to rank documents using document identifiers (docIDs). As is standard, we assume the ARR model uses token embeddings $\operatorname{E}\in\mathbb{R}^{|\mathscr{V}|\times n}$ and computes next-token probabilities over $\mathscr{V}$ as $P(\cdot\mid c)=\sigma(\operatorname{E}\varphi(c))$ where $\varphi(c)\in\mathbb{R}^{n}$ is a hidden vector produced by the ARR model given context $c$ and $\sigma$ is the softmax function.
In the setting of document ranking, documents can be represented by unique identifiers, or docIDs, where each docID is a sequence of tokens from $\mathscr{V}_{\text{docIDs}}\subseteq\mathscr{V}$. To ensure the model only generates valid docID tokens we can make use of constrained decoding, which restricts next-token probabilities to those tokens in $\mathscr{V}_{\text{docIDs}}$ by computing

### v2

@@ Abstract
The success of Large Language Models (LLMs) has motivated a shift toward generative approaches to retrieval and ranking, aiming to supersede classical Dual Encoders (DEs) and Cross Encoders (CEs). A prominent paradigm is pointwise Autoregressive Ranking (ARR), where an LLM generates document identifiers (docIDs) token-by-token to enable ranking via beam search. ARR offers the promise of superior expressivity compared to DEs while avoiding the prohibitive computational cost of CEs. However, a formal theoretical foundation for this expressive power has been missing. Moreover, the standard next-token prediction loss is rank-agnostic and inappropriate for finetuning an LLM for ranking tasks.
In this paper, we first prove that the expressive capacity of ARR is strictly superior to DEs. While a DE requires an embedding dimension that grows linearly with corpus size to achieve arbitrary rankings, ARR can solve it with a constant hidden dimension. We then propose SToICaL (Simple Token-Item Calibrated Loss), a generalized rank-aware training loss for LLM finetuning. By using item-level reweighting and prefix-tree marginalization, we distribute probability mass over valid docID tokens based on their ground-truth relevance. Experiments on WordNet and ESCI datasets verify that our loss suppresses invalid docID generations and significantly improves ranking metrics beyond top-1 retrieval.
@@ 1 Introduction
Ranking documents by their relevance to a user query is a cornerstone of modern information retrieval. The dominant paradigm for this task is a multi-stage pipeline. The first stage, retrieval, aims to efficiently select a candidate set of documents from a massive corpus. This is typically accomplished using Dual Encoder (DE) models, which represent queries and documents as dense vectors and use fast approximate nearest neighbor (ANN) search to find the top matches (karpukhin2020dense; guo2020accelerating). While efficient, the expressive power of DEs is inherently limited, as the rich interaction between a query and a document is compressed into a single vector similarity. Consequently, a second reranking stage is often required. In this stage, a more powerful but computationally expensive Cross Encoder (CE) model jointly processes the query and each candidate document to produce a more accurate relevance score (nogueira2019passage). CEs are infeasible for first-stage retrieval due to their prohibitive cost, which scales linearly with corpus size.
@@
The success of LLMs has inspired a new, unified approach to retrieval and ranking in lieu of the two-stage approach (metzler2021rethinking; li2023large; lee2024can). In this paradigm, an LLM is trained to function as a generative retrieval model, autoregressively decoding the relevant documents or their identifiers (docIDs) token-by-token, conditioned on the query (tay2022transformer). For ranking, this model can generate a sorted list of documents by employing beam search, effectively ranking documents by their conditional generation probability. This approach, which we dub Autoregressive Ranking (ARR), offers several compelling advantages over the traditional DE/CE pipeline. It collapses the two-stage process into a single model, eliminating the need for a separate ANN index. Furthermore, it is more efficient than CEs as it does not require scoring every document individually, while preserving CEs’ cross-attention over already generated docID tokens. Crucially, it has also been argued that by generating documents token-by-token, ARR models can capture deeper and more nuanced query-document interactions than DEs (yuan2024generative); while intuitive, the evidence for this has been primarily empirical. In this paper, we address two critical gaps in the understanding and optimization of ARR models for ranking:
@@
  • We provide a theoretical foundation for the superior expressive capacity of ARR over DEs. A rigorous analysis of the embedding geometry required for ranking shows that for a DE to achieve any ordering of $k$ documents, its embedding dimension must grow linearly with $k$. In contrast, we prove that an ARR model with constant hidden dimension is theoretically sufficient to rank an arbitrary number of documents. This offers a formal explanation for the advantages of ARR.
@@
  • We introduce SToICaL, a rank-aware generalization of next-token prediction loss for enhanced ranking capabilities. Most existing generative retrieval methods are trained with standard next-token prediction, an objective which fails to explicitly model the relative ordering of documents, which is crucial for ranking. We introduce a new loss function designed to make the LLM aware of the entire ranked list during training, while keeping the generation and signal length on the order of a single docID rather than a full list, enabling ranking via beam search. We test our method on WordNet and a shopping queries dataset, demonstrating its superior ranking performance.
@@ 2 Related Work
Classical IR Losses
Training objectives for retrieval and ranking can be broadly grouped into the categories of pointwise, pairwise, and listwise (liu2009learning). A pointwise loss as in chen2009ranking of the form $\mathcal{L}(q,d)$ encourages the model to assign high scores to more relevant items and vice versa; each training example consists of a single document to penalize independently of the rest. Pairwise loss functions (cao2007learning) of the form $\mathcal{L}(q,d^{+},d^{-})$ strive to maximize the margin between relevant and irrelevant documents. Listwise losses such as xia2008listwise of the form $\mathcal{L}(q,d_{1},\ldots,d_{n_{q}})$ jointly take as input the model scores of all documents under consideration for a given query, and can encourage the most flexible scoring configurations at the granularity of permutations rather than individual items and commonly rely on the Plackett-Luce probability model over permutations (luce1959individual; plackett1975analysis).
@@
Autoregressive decoder LLMs introduce the facet of text generation on top of the simple API of “model-as-document-scorer”. The score-based loss formulations discussed above remain applicable insofar as a docID’s probability can be obtained from the probabilities of its constituent tokens; notable LLM-based reranking methods (sachan2022improving; drozdov2023parade) rely specifically on LLM-based log-likelihood to obtain scores. Meanwhile, the terminology of “pointwise” vs “listwise” naturally acquires a second meaning, i.e., whether the text string which the LLM generates comprises a single docID or a ranked list of docIDs (separated by a delimiter and in descending order, as per common convention). Our “pointwise” approach relates to both these senses (scoring just one document, or generating just a single docID string), since steering LLM probabilities to be more rank-aware also directly impacts the rankings produced by a decoding strategy like beam search at inference time. The generative approach naturally allows us to perform ranking with a single model call, in contrast to CEs which evaluate the score for every document separately. The compact pointwise output format mitigates snowballing errors and is efficient to train under our rank-aware modification to next-token prediction.
@@
A handful of examples illustrate the diverse reranking paradigms afforded by generative LLMs. RankVicuna (pradeep2023rankvicuna) and RankZephyr (pradeep2023rankzephyr) explore training/distillation choices, using instruction tuning with next token prediction, to achieve a robust open-source zero-shot LLM for listwise ranking. LTRGR (li2024learning), on top of a next-token-prediction-based “learning-to-generate” phase, introduces a “learning-to-rank” phase that maximizes the margin between the autoregressively-computed scores of relevant and irrelevant passage identifiers. RecRanker (luo2024recranker) trains an LLM with instruction tuning on a combination of pointwise, pairwise, and listwise ranking objectives for top-$k$ sequential recommendation, and at inference time aggregates the three types of responses into a hybrid ranking by means of respective utility functions. GenRank (huang2025towards) identify the autoregressive generation mechanism as critical to the effectiveness of generative ranking, more so than the pre-training paradigm. sun2023chatgpt explore the ranking capabilities of ChatGPT and GPT-4 and propose a distillation technique to learn smaller specialized rankers. gupta2025scalable introduce BlockRank, efficiently tailoring the attention mechanism to in-context ranking. We argue that autoregressive methods which employ next-token prediction or instruction fine-tuning are relying on a token-level loss which is fundamentally not rank-aware. Our proposed loss function is a simple, principled and effective injection of rank-awareness into next-token prediction — a loss function which is ubiquitous to autoregressive models.
@@
Differentiable Search Index (DSI) (tay2022transformer) instills the docID index in model weights as part of the training objective. A critical component of parametric generative retrieval is effective docID design (li2023multiview) — we take up this question for the ESCI dataset in Section˜5.1. Note that we opt to use in-context ranking (gupta2025scalable), which leverages the LLM’s in-context reasoning and copying abilities to perform retrieval and ranking from a pool of documents, with the docIDs provided as part of the prompt wang2024large. For in-context ranking, one can interpret the longer prompt as an “augmented query”; our theory (Section˜3) and loss function (Section˜4), however, generalizes to parametric as well as in-context ranking.
@@ 3 The Capacity for Ranking
Let $\mathcal{Q}$ be a set of queries and $\mathcal{D}$ be a corpus of documents (or their docIDs). Towards establishing a theory of the ARR’s model capacity for ranking, we start with a formal definition of the ranking task as follows.
@@
We consider a ranking task where for each query $q\in\mathcal{Q}$, the goal is to generate a list $\mathscr{L}(q)=[d_{1}(q),\ldots,d_{k}(q)]$ of relevant documents, $d_{i}(q)\in\mathcal{D}$. Ordering is significant: $\forall i<j$, $d_{i}(q)$ is considered more relevant to $q$ than $d_{j}(q)$, denoted by $d_{i}(q)\succ d_{j}(q)$.
We refer to Autoregressive Ranking (ARR) as the use of a causal language model for ranking tasks. Such a model predicts the probability of a sequence of tokens conditioned on its preceding context. Given a query $q$, an ARR model defines a conditional probability distribution over document token sequences. The probability of any document $d\in\mathcal{D}$ conditioned on a query $q\in\mathcal{Q}$ is the product of $d$’s token probabilities, calculated autoregressively. [^2] Fortunately, the autoregressive nature of causal models offers a more efficient solution: Instead of scoring, we can directly generate the high-probability documents using greedy decoding or beam search. By employing beam search, we can generate the most likely documents conditioned on the query, directly yielding a ranked list without the need for exhaustive computation.
Both ARR and dual encoders (DEs) are examples of a scoring-based ranking architecture. Such architectures generate a score for each document given a query $q$, with the objective that the scores, when sorted in descending order, align with the ground-truth ordering of the relevant documents, and that scores for irrelevant documents are smaller than scores for relevant ones. We state this formally below.
@@
A (parametric) scoring-based ranking architecture is a function $f(q,d;\theta):\mathcal{Q}\times\mathcal{D}\rightarrow\mathbb{R}$, which takes a pair $q,d$ and produces a scalar score interpreted as the similarity. Here, $\theta$ denotes learnable model parameters.
In particular, we say that a scoring-based architecture $f(q,d;\theta)$ can solve a ranking task if there exists a $\theta$ such that the scores solve the ranking task for all queries in $\mathcal{Q}$.
To simplify our analysis, we consider an idealized case of the ranking task defined as follows.
@@
Given $\mathcal{Q}$ and $\mathcal{D}$, we call a ranking task complete if:
  • $\forall q\in\mathcal{Q}$, all documents are relevant, i.e., $\mathcal{D}(q)=\mathcal{D}$
  • $\forall$ ordering $\pi(\mathcal{D})$ of the documents $\mathcal{D}$, $\exists q\in\mathcal{Q}$ s.t. its ranked list of relevant documents $\mathscr{L}(q)\equiv\pi(\mathcal{D})$.
We proceed to show that the architectural capacity of ARR for ranking is not limited in the same way as that of a DE.
@@ 3.1 Ranking via Dual Encoders
In this section, we start with an analysis of a DE’s capacity for the ranking task. Our result shows that an embedding dimension that grows linearly with the number of relevant documents is necessary for a DE to correctly rank them.
In DEs, the query and document are embedded into the same Euclidean space, with the negated Euclidean distance as the similarity score. Specifically, a dual encoder generates a similarity score for each pair of $q\in\mathcal{Q}$ and $d\in\mathcal{D}$ as
$f_{\text{DE}}(q,d;\theta)=-\|E_{Q}(q;\theta_{\mathcal{Q}})-E_{D}(d;\theta_{\mathcal{D}})\|_{2}$ (1)
where $E_{Q}:\mathcal{Q}\to\mathbb{R}^{n}$ and $E_{D}:\mathcal{D}\to\mathbb{R}^{n}$ are query and document encoders, respectively, and $n$ is the dimension of the embedding space. Here, we assume that $E_{Q}$ and $E_{D}$ can have an arbitrary architecture. This means that they can be taken to be as expressive as needed when studying whether or not DEs are expressive enough to solve the ranking task. Our main result for DEs is the following.
Proposition 3.1 (DE Insufficiency for Complete Ranking).
Let $f_{\text{DE}}(q,d;\theta)$ be a dual encoder architecture defined in Equation˜1 with some embedding dimension $n$, and let $k:=|\mathcal{D}|$. Then there does not exist a $\theta$ such that $f_{\text{DE}}(q,d;\theta)$ solves a complete ranking task when
@@
Let $N_{n,p}(k)$ denote the maximum number of distinct distance permutations generated by $k$ sites in $\mathbb{R}^{n}$ with the $L_{p}$ metric. From skala2009counting Corollary 8, $N_{n,2}(k)\leq k^{2n}$. When Inequality (2) holds, we have $k^{2n}<k^{!}$, and thus $n$-dimensional embeddings are unable to capture all $k!$ permutations required in a complete ranking task. ∎
Using Stirling’s formula which states $\ln(k!)=k\ln k-k+O(\ln(k))$, the condition in Equation˜2 becomes $n<\frac{k}{2}-\frac{k}{2\ln k}+O(1)$. Since $k/\ln k$ grows sublinearly in $k$, this result states that the embedding dimension $n$ of the DE needs to grow linearly with the number of documents $k$.
@@
3.2 Ranking via (Pointwise) Autoregressive Rankers
We turn our attention to the ability of an ARR to rank documents using docIDs. [^3] Proofs for this section can be found in Appendix A. As is standard, we assume the ARR uses token embeddings $\operatorname{E}\in\mathbb{R}^{|\mathcal{V}|\times n}$ and computes next-token probabilities over the vocabulary $\mathcal{V}$ as $P(\cdot\mid c)=\sigma(\operatorname{E}\varphi(c))$ where $\varphi(c)\in\mathbb{R}^{n}$ is a hidden vector produced by the ARR given context $c$ and $\sigma$ is the softmax.
In the setting of document ranking, documents can be represented by unique identifiers, or docIDs, where each docID is a sequence of tokens from $\mathscr{V}_{\text{docIDs}}\subseteq\mathcal{V}$. To ensure the model only generates valid docID tokens we can make use of constrained decoding, which restricts next-token probabilities to those tokens in $\mathscr{V}_{\text{docIDs}}$ by computing